Calcium hydroxide and sulfuric acid react to produce calcium sulfate and water according to the equation:

Ca(OH)2(aq)  +  H2SO4(aq)  --> CaSO4(s)  +  2 H2O(l)

How many liters of 0.8 M Ca(OH)2 are needed to produce 4.4 moles of CaSO4?



1
Expert's answer
2021-04-13T08:04:12-0400

Molar Mass of Ca(OH)2 = 74.1 g/mol

Molar Mass of CaSO4 = 136.14

Ratio = 1:1

4.4 moles of CaSO4 = 4.4 moles of Ca(OH)2

= 74.1 × 4.4

= 326.04 × 0.8 = 260.832 g

= 0.260832 L